OpenStudy (anonymous):

Use the fundamental identities to simplify the expression. csc theta/sec theta

OpenStudy (lgbasallote):

\[\large \frac{\frac{1}{\sin \theta}}{\frac{1}{\cos \theta}}\] that give you any ideas?

OpenStudy (anonymous):

yes! thank you!!